Chevy Volt main battery has capacity to store 16.5 kWh of electrical energy with the nominal voltage equal to 390V If we would like to store this amount of energy in a capacitor with the same voltage, then what would have to be the capacitance of that capacitor?
Answer #1

U = 0.5*Q*V


16.5*3.6*10^6 = 0.5*Q*390


Q = 59.4*10^6/(0.5*390) = 3.046*10^5 C